# Bi-Mamba4TS: Bidirectional Mamba for Time Series Forecasting

### v1

@@ Introduction
Weather 96 0.164 0.212 0.170 0.214 0.174 0.214 0.178 0.219 0.162 0.232 0.266 0.336 0.196 0.235 0.172 0.220 0.186 0.237 0.178 0.223
192 0.214 0.256 0.220 0.257 0.221 0.254 0.224 0.259 0.208 0.251 0.307 0.367 0.241 0.271 0.219 0.261 0.233 0.273 0.223 0.261
336 0.269 0.296 0.276 0.298 0.278 0.298 0.292 0.306 0.270 0.305 0.359 0.395 0.292 0.306 0.280 0.306 0.289 0.312 0.288 0.309
720 0.348 0.349 0.356 0.349 0.358 0.349 0.354 0.348 0.398 0.418 0.419 0.428 0.363 0.353 0.365 0.359 0.356 0.352 0.372 0.363
Traffic 96 0.383 0.261 0.388 0.267 0.395 0.268 0.457 0.295 0.534 0.280 0.613 0.388 0.647 0.384 0.593 0.321 0.676 0.407 1.037 0.441
192 0.398 0.269 0.411 0.274 0.417 0.276 0.471 0.299 0.556 0.300 0.616 0.382 0.596 0.359 0.617 0.336 0.631 0.386 1.061 0.455
336 0.410 0.276 0.423 0.280 0.433 0.283 0.482 0.304 0.578 0.315 0.616 0.382 0.601 0.361 0.629 0.336 0.640 0.387 1.095 0.470
720 0.447 0.298 0.458 0.300 0.467 0.302 0.514 0.322 0.594 0.321 0.622 0.337 0.642 0.381 0.640 0.350 0.681 0.402 1.121 0.474
Electricity 96 0.142 0.238 0.142 0.238 0.148 0.240 0.174 0.259 0.152 0.249 0.201 0.317 0.206 0.288 0.168 0.272 0.217 0.304 0.237 0.335
192 0.157 0.253 0.169 0.267 0.162 0.253 0.178 0.265 0.174 0.275 0.222 0.334 0.206 0.290 0.184 0.289 0.216 0.306 0.258 0.350
336 0.172 0.271 0.178 0.275 0.178 0.269 0.196 0.282 0.197 0.294 0.231 0.338 0.220 0.305 0.198 0.300 0.232 0.321 0.273 0.362
720 0.200 0.297 0.207 0.303 0.225 0.317 0.237 0.316 0.258 0.339 0.254 0.361 0.252 0.337 0.220 0.320 0.273 0.352 0.300 0.382
ETTh1 96 0.376 0.403 0.386 0.406 0.386 0.405 0.393 0.408 0.395 0.411 0.449 0.459 0.387 0.406 0.384 0.402 0.382 0.398 0.414 0.419
192 0.411 0.425 0.439 0.437 0.441 0.436 0.445 0.434 0.460 0.447 0.500 0.482 0.439 0.435 0.436 0.429 0.427 0.425 0.464 0.448
336 0.455 0.445 0.485 0.467 0.487 0.458 0.474 0.451 0.479 0.455 0.521 0.496 0.493 0.457 0.491 0.469 0.486 0.487 0.516 0.478
720 0.460 0.464 0.495 0.490 0.503 0.491 0.480 0.471 0.480 0.469 0.514 0.512 0.490 0.478 0.521 0.500 0.472 0.468 0.538 0.509
ETTh2 96 0.289 0.341 0.298 0.349 0.297 0.349 0.302 0.348 0.311 0.372 0.346 0.388 0.305 0.352 0.340 0.374 0.302 0.349 0.325 0.364
192 0.367 0.389 0.386 0.401 0.380 0.400 0.388 0.400 0.393 0.414 0.456 0.452 0.424 0.439 0.402 0.414 0.382 0.400 0.433 0.427
336 0.410 0.424 0.412 0.425 0.428 0.432 0.426 0.433 0.437 0.449 0.482 0.486 0.456 0.473 0.452 0.452 0.421 0.439 0.471 0.457
720 0.421 0.439 0.424 0.445 0.428 0.432 0.426 0.433 0.437 0.449 0.482 0.486 0.476 0.493 0.462 0.468 0.437 0.458 0.499 0.480
ETTm1 96 0.312 0.354 0.329 0.364 0.334 0.368 0.329 0.367 0.327 0.363 0.505 0.475 0.353 0.374 0.338 0.375 0.340 0.374 0.375 0.402
192 0.358 0.383 0.371 0.387 0.377 0.391 0.367 0.385 0.376 0.397 0.553 0.496 0.389 0.391 0.374 0.387 0.377 0.390 0.427 0.434
336 0.388 0.404 0.403 0.409 0.426 0.420 0.399 0.419 0.424 0.425 0.621 0.537 0.421 0.413 0.410 0.411 0.401 0.407 0.455 0.452
720 0.444 0.433 0.471 0.448 0.491 0.459 0.454 0.439 0.482 0.462 0.671 0.561 0.484 0.448 0.478 0.450 0.453 0.442 0.527 0.488
ETTm2 96 0.174 0.259 0.181 0.264 0.180 0.264 0.175 0.259 0.181 0.266 0.255 0.339 0.182 0.264 0.187 0.267 0.177 0.261 0.191 0.272
192 0.240 0.304 0.251 0.311 0.250 0.309 0.241 0.302 0.244 0.307 0.281 0.340 0.257 0.315 0.249 0.309 0.240 0.298 0.261 0.316
336 0.303 0.345 0.311 0.348 0.311 0.348 0.305 0.343 0.310 0.351 0.339 0.372 0.318 0.353 0.321 0.351 0.305 0.345 0.330 0.358
720 0.402 0.407 0.410 0.404 0.412 0.407 0.402 0.400 0.406 0.404 0.433 0.432 0.426 0.419 0.408 0.403 0.403 0.400 0.450 0.427
@@
  • CrossGNN (Huang et al. 2024) is a model mainly based on GNN. It views time series in a multi-scale way. It uses GNN to capture both intra-scale and inter-series dependencies.
  • D-Mamba (Wang et al. 2024) generates embeddings for each time series through a simple MLP layer and uses Mamba to extract inter-series dependencies.
@@
We set the same historical look-back window size of $L=96$ for all models on all datasets. The predicting length is set to $H\in\{96,192,336,720\}$. We set $P=24$ and $S=12$ on all datasets. Considering that different models may have different sensitivity of granularity of local information, we use the patching settings consistent with the original papers for PatchTST and Crossformer. For the SRA decider, we set $\lambda=0.6$ because generally when the Pearson coefficient correlation is larger than $0.6$, it can be inferred that the two sequences have a strong correlation. More detailed evaluation of $\lambda$ is shown in Fig. 3. As for the Mamba block, we set $d\_conv=2$ and $expand=1$ on all datasets. Considering that datasets with more variables may have more complex patterns, we set $d\_model=64$, $d\_state=32$ for Weather, Electricity and Traffic and $d\_model=32$, $d\_state=2$ for ETT datasets. We evaluate Bi-Mamba4TS with encoder layer $l\in\{1,2,3\}$ and conduct grid search for learning rate on [4e-5, 1e-4, 4e-4, 1e-3, 4e-3, 1e-2] to find the best results for all datasets. For the metrics to evaluate the performance, we utilize MSE and MAE, defined in Eq 5 and Eq 6. All models are optimized using the ADAM algorithm(Kingma and Ba 2014) and early stopping mechanism. The training process is limited to 60 epochs.
$\mathrm{MSE}(Y,\hat{Y})$ $=\frac{1}{|Y|}\sum_{i=1}^{|Y|}|Y_{i}-\hat{Y}_{i}|^{2}$ (5)
$\mathrm{MAE}(Y,\hat{Y})$ $=\frac{1}{|Y|}\sum_{i=1}^{|Y|}|Y_{i}-\hat{Y}_{i}|$ (6)
Table 4: Ablation studies of (a) Bi-Mamba4TS-I which use channel-independent strategy only, (b) Bi-Mamba4TS-M which use channel-mixing strategy only, (c) Mamba4TS which use forward direction Mamba block only and (d) D-Mamba and (d) PatchTST used for the benchmark models.
Models Bi-Mamba4TS-I Bi-Mamba4TS-M Mamba4TS D-Mamba PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity 96 0.163 0.250 0.142 0.238 0.149 0.245 0.142 0.238 0.174 0.259
192 0.171 0.259 0.157 0.253 0.163 0.258 0.169 0.267 0.178 0.265
336 0.188 0.276 0.172 0.271 0.182 0.278 0.178 0.275 0.196 0.282
720 0.229 0.311 0.200 0.297 0.208 0.302 0.207 0.303 0.237 0.316
ETTh1 96 0.376 0.403 0.394 0.410 0.383 0.403 0.386 0.406 0.393 0.408
192 0.411 0.425 0.450 0.439 0.431 0.430 0.439 0.437 0.445 0.434
336 0.455 0.445 0.492 0.454 0.474 0.450 0.485 0.467 0.474 0.451
720 0.460 0.464 0.511 0.490 0.462 0.465 0.495 0.490 0.480 0.471
@@
We conduct all experiments on the same device. The detailed information about our experimental platform is shown in table 2.
@@
Table 3 shows the results of long-term multivariate forecasting. Our model achieves superior performance compared with other baselines on all datasets with different predicting lengths. Compared with the current SOTA Transformer model iTransformer, Bi-Mamba4TS reduces the MSE, MAE errors by 4.92% and 2.16% on average. The improvement of MSE and MAE comes to 3.55% and 1.87% compared with D-Mamba, which is also an SSM-based model using Mamba. Note that Bi-Mamba4TS adopts channel-mixing strategy on Weather, Traffic and Electricity and channel-independent strategy on 4 ETT datasets. For channel-mixing strategy, the patching approach used by Bi-Mamba4TS helps the model capture inter-series dependencies at a finer granularity, therefore achieving more accurate predicting. For channel-independent strategy, the patch-wise tokens fed into Mamba blocks contain more semantic information, enhancing the model’s ability for capturing the inner evolutionary process of each univariate time series. This also indicates that adopting different tokenization strategies for specific datasets help improve the model’s performance.
Figure 3: The tokenization strategy indicator $ts$ of different datasets with varying settings of filter threshold $\lambda$.
@@
We verify the effectiveness of the SRA decider and the bidirectional Mamba settings. Before we conduct the ablation experiments, we first export the results of the SRA decider to view the strategy choosing for different datasets. For Pearson correlation coefficient $\rho$, we typically use $\lambda>0.6$ to evaluate whether two variables have positive correlation strong enough to influence each other. We choose $\lambda=0.6$ and calculate the strategy indicator $ts$ for different datasets, the results are shown in Fig. 3. For ETT datasets that have few variables, the model tends to adopt channel-independent strategy to emphasis intra-series dependencies, while for Weather, Traffic and Electricity that have many variables, the model tends to adopt channel-mixing strategy to emphasis inter-series dependencies. We remove the SRA decider to form two specific Bi-Mamba4TS models: (a) Bi-Mamba4TS-I and (b) Bi-Mamba4TS-M that only use channel-independent or channel-mixing strategy respectively. For the bidirectional Mamba blocks, we remove the backward direction Mamba to evaluate the effectiveness of bidirectional design. This setting forms the following model: (c) Mamba4TS. We choose the results for Electricity and ETTh1, as shown in table 4. Without the decider, we mainly focus on the results of Bi-Mamba4TS-I on Electricity and Bi-Mamba4TS-M on ETTh1. The MSE and MAE increase by 10.65% and 3.35% on Electricity and 7.86% and 3.15% on ETTh1 respectively. Therefore, utilizing different tokenization strategies on different datasets brings better model performance. In addition, the proposed SRA decider makes the optimal choice of channel-independent and channel-mixing strategies. It is worth noting that Bi-Mamba4TS-I shows better performance on all cases than PatchTST that uses patching technique and channel-independent tokenization strategy. This indicates that Mamba demonstrates superior capabilities in long sequence modeling beyond self-attention. When the backward Mamba block is removed, the predicting performance is also deteriorated to some extent. This indicates that the bidirectional design helps the model get more comprehensive understanding of the MTS data.
@@
To verify whether Bi-Mamba4TS is sensitive to the hyper-parameters, we conduct experiments on (a) Pearson coefficient filter $\lambda$, (b) dropout, (c) Conv-1D kernel size of Mamba and (d) state dimension size. For each setting, we repeat the experiment 5 times with 60 epochs (with early-stop control) and record the average MSE of $H\in\{96,192,336,720\}$. Overall, different parameter settings do not cause significant changes in the results, which also shows the robustness of our proposed model.
Pearson coefficient filter $\lambda$.
The value $\lambda$ is used for filter out variable pairs that have positive correlations. A larger $\lambda$ corresponds to stronger positive correlations between two sequences. We set $\lambda\in\{0.2,0.4,0.6,0.8\}$ and calculate the decision ratio $r$ and decision indicator $ts$. The results are shown in Fig. 3. Overall, different $\lambda$ does not frequently cause the change of tokenization strategy. For datasets that have large number of variables, such as Traffic and Electricity, the decision indicator $ts$ prefers to be equal to $1$, while for datasets with less variables such as ETTh1 and ETTh2, $ts$ is more likely to be 00. Note that the $ts$ may undergo a change from $\lambda=0.2$ to $\lambda=0.4$ on Weather and ETTh2, resulting in different tokenization strategies. However, a Pearson coefficient closed to $0.2$ can only indicate a weak relationship between the two sequences, therefore, this threshold is not a general option. We mainly focus on the decision performance of $\lambda\geq 0.4$ and the results show that the strategy decision is stable.
Dropouts.
@@
The 1-D convolution module in Mamba is used for capturing temporal dependencies of the input series. We conduct experiments on different kernel size of 1, 2 and 3. The results in Fig. LABEL:fig:hyper(e) and Fig. LABEL:fig:hyper(f) show that Bi-Mamba4TS gets optimal performance with the kernel size of 2. Overall, our model is not sensitive to the kernel size.
@@
We conduct the following experiments to comprehensively evaluate the model efficiency from (a) forecasting performance, (b) memory footprint and (c) training speed. We set $L=96,H=192$ as the forecasting task and use $Batch=32$ for ETTh2 and $Batch=16$ for Traffic. We train all the models according to their optimal settings that achieve the best performance. The results are shown in Fig. LABEL:fig:efficiency(a) and Fig. LABEL:fig:efficiency(b). Our proposed Bi-Mamba4TS strikes a good balance among predicting performance, training speed and memory usage. Overall, Mamba’s computational cost is comparable to the self-attention mechanism in Transformer, which is reflected in the training speed. However, it is worth noting that with the same number of model layers, a model using Mamba costs less memory than that using a Transformer-based model. With the better predicting performance, faster training speed and less memory costs, Mamba shows the great potential in LTSF.

### v3

@@ Introduction
Weather 96 0.159 0.205 0.166 0.210 0.174 0.214 0.178 0.219 0.158 0.230 0.266 0.336 0.196 0.235 0.172 0.220 0.186 0.237 0.178 0.223
192 0.205 0.249 0.215 0.253 0.221 0.254 0.224 0.259 0.206 0.277 0.307 0.367 0.241 0.271 0.219 0.261 0.233 0.273 0.223 0.261
336 0.264 0.291 0.276 0.298 0.278 0.298 0.292 0.306 0.272 0.335 0.359 0.395 0.292 0.306 0.280 0.306 0.289 0.312 0.288 0.309
720 0.343 0.344 0.353 0.349 0.358 0.349 0.354 0.348 0.398 0.418 0.419 0.428 0.363 0.353 0.365 0.359 0.356 0.352 0.372 0.363
Traffic 96 0.375 0.258 0.381 0.261 0.395 0.268 0.457 0.295 0.522 0.290 0.613 0.388 0.647 0.384 0.593 0.321 0.676 0.407 1.037 0.441
192 0.394 0.269 0.397 0.267 0.417 0.276 0.471 0.299 0.530 0.293 0.616 0.382 0.596 0.359 0.617 0.336 0.631 0.386 1.061 0.455
336 0.406 0.274 0.423 0.276 0.433 0.283 0.482 0.304 0.558 0.305 0.616 0.382 0.601 0.361 0.629 0.336 0.640 0.387 1.095 0.470
720 0.440 0.288 0.458 0.300 0.467 0.302 0.514 0.322 0.589 0.328 0.622 0.337 0.642 0.381 0.640 0.350 0.681 0.402 1.121 0.474
Electricity 96 0.140 0.238 0.142 0.238 0.148 0.240 0.174 0.259 0.219 0.314 0.201 0.317 0.206 0.288 0.168 0.272 0.217 0.304 0.237 0.335
192 0.155 0.253 0.169 0.267 0.162 0.253 0.178 0.265 0.231 0.322 0.222 0.334 0.206 0.290 0.184 0.289 0.216 0.306 0.258 0.350
336 0.170 0.269 0.178 0.275 0.178 0.269 0.196 0.282 0.246 0.337 0.231 0.338 0.220 0.305 0.198 0.300 0.232 0.321 0.273 0.362
720 0.197 0.293 0.207 0.303 0.225 0.317 0.237 0.316 0.280 0.363 0.254 0.361 0.252 0.337 0.220 0.320 0.273 0.352 0.300 0.382
ETTh1 96 0.378 0.395 0.386 0.406 0.386 0.405 0.393 0.408 0.423 0.448 0.449 0.459 0.387 0.406 0.384 0.402 0.382 0.398 0.414 0.419
192 0.427 0.428 0.448 0.444 0.441 0.436 0.445 0.434 0.471 0.474 0.500 0.482 0.439 0.435 0.436 0.429 0.427 0.425 0.464 0.448
336 0.471 0.445 0.494 0.468 0.487 0.458 0.474 0.451 0.570 0.546 0.521 0.496 0.493 0.457 0.491 0.469 0.486 0.487 0.516 0.478
720 0.470 0.457 0.493 0.488 0.503 0.491 0.480 0.471 0.653 0.621 0.514 0.512 0.490 0.478 0.521 0.500 0.482 0.477 0.538 0.509
ETTh2 96 0.291 0.342 0.298 0.349 0.297 0.349 0.302 0.348 0.745 0.584 0.346 0.388 0.305 0.352 0.340 0.374 0.302 0.349 0.325 0.364
192 0.368 0.392 0.379 0.398 0.380 0.400 0.388 0.400 0.877 0.656 0.456 0.452 0.424 0.439 0.402 0.414 0.382 0.400 0.433 0.427
336 0.407 0.424 0.417 0.432 0.428 0.432 0.426 0.433 1.043 0.731 0.482 0.486 0.456 0.473 0.452 0.452 0.421 0.439 0.471 0.457
720 0.421 0.439 0.431 0.449 0.428 0.432 0.431 0.446 1.104 0.763 0.515 0.511 0.476 0.493 0.462 0.468 0.437 0.458 0.499 0.480
ETTm1 96 0.320 0.360 0.331 0.368 0.334 0.368 0.329 0.367 0.404 0.426 0.505 0.475 0.353 0.374 0.338 0.375 0.340 0.374 0.375 0.402
192 0.361 0.383 0.371 0.387 0.377 0.391 0.367 0.385 0.450 0.451 0.553 0.496 0.389 0.391 0.374 0.387 0.377 0.390 0.427 0.434
336 0.386 0.402 0.417 0.418 0.426 0.420 0.399 0.419 0.532 0.515 0.621 0.537 0.421 0.413 0.410 0.411 0.401 0.407 0.455 0.452
720 0.445 0.437 0.471 0.448 0.491 0.459 0.454 0.439 0.666 0.589 0.671 0.561 0.484 0.448 0.478 0.450 0.453 0.442 0.527 0.488
ETTm2 96 0.176 0.263 0.179 0.263 0.180 0.264 0.175 0.259 0.287 0.366 0.255 0.339 0.182 0.264 0.187 0.267 0.177 0.261 0.191 0.272
192 0.242 0.304 0.253 0.310 0.250 0.309 0.241 0.302 0.414 0.492 0.281 0.340 0.257 0.315 0.249 0.309 0.240 0.298 0.261 0.316
336 0.304 0.344 0.312 0.348 0.311 0.348 0.305 0.343 0.597 0.542 0.339 0.372 0.318 0.353 0.321 0.351 0.305 0.345 0.330 0.358
720 0.402 0.402 0.412 0.408 0.412 0.407 0.402 0.400 1.730 1.042 0.433 0.432 0.426 0.419 0.408 0.403 0.403 0.400 0.450 0.427
Solar 96 0.184 0.222 0.205 0.241 0.203 0.237 0.234 0.286 0.310 0.331 0.884 0.711 0.290 0.378 0.250 0.292 0.222 0.314 0.238 0.285
192 0.225 0.254 0.237 0.270 0.233 0.261 0.267 0.310 0.734 0.725 0.834 0.692 0.320 0.398 0.296 0.318 0.240 0.318 0.282 0.302
336 0.249 0.270 0.256 0.286 0.248 0.273 0.290 0.315 0.750 0.735 0.941 0.723 0.353 0.415 0.319 0.330 0.264 0.324 0.329 0.345
720 0.250 0.273 0.260 0.288 0.249 0.275 0.289 0.317 0.769 0.765 0.882 0.717 0.356 0.413 0.338 0.337 0.260 0.321 0.326 0.348
@@
  • iTransformer (Liu et al. 2023) inverts the modeling method of Transformer by using self-attention to capture inter-series dependencies.
@@
  • CrossGNN (Huang et al. 2024) is a model mainly based on GNN. It views time series in a multi-scale way. It uses GNN to capture both cross-scale and cross-series dependencies.
  • S-Mamba (Wang et al. 2024) generates embeddings for each time series through a simple MLP layer and uses Mamba to extract inter-series dependencies.
@@
We set the same historical look-back window size of $L=96$ for all models on all datasets. The predicting length is set to $H\in\{96,192,336,720\}$. We set $S=\frac{1}{2}P$ and use patch length $P=\frac{1}{4}L$ by default. Considering that different models may have different sensitivity of granularity of local information, we maintain the patching settings consistent with the original papers for PatchTST and Crossformer. For the SRA decider, we set $\lambda=0.6$ because generally when the Spearman coefficient correlation is larger than $0.6$, it can be inferred that the two sequences have a strong positive correlation. More detailed evaluation of $\lambda$ is shown in Fig. LABEL:fig:decision_maker. Bi-Mamba+ splits a sequence into patches, which mainly retain local information but contains less global information compared to a whole sequence. Considering that datasets with more variables may contain more complex evolutionary patterns, for Bi-Mamba+, PatchTST and Crossformer that use patching technique, we set $D=128$ for Weather, Traffic, Electricity, Solar and $D=64$ for ETT datasets, while for S-Mamba and iTransformer that map the whole sequence to tokens, we set $D=512$ for Weather, Traffic, Electricity, Solar and $D=256$ for ETT datasets. As for parameters within Mamba+ block, we follow (Ahamed and Cheng 2024; Wang et al. 2024) and set convolutional kernel size $d\_conv=2$ and hidden state expansion $expand=1$ on all datasets. We set hidden dimension $d\_state=16$ for Weather, Electricity and Traffic and $d\_state=8$ for ETT datasets. We evaluate Bi-Mamba+ with encoder layer $l\in\{1,2,3\}$ and conduct grid search for learning rate on [5e-5, 1e-4, 2e-4, 5e-4, 1e-3, 2e-3, 5e-3] to find the best results. For the metrics to evaluate the performance, we utilize MSE and MAE. All models are optimized using the ADAM algorithm(Kingma and Ba 2014) and early stopping mechanism. The training process is limited to 40 epochs.
@@
To exclude the influence of the model training environment, we conduct all experiments on the same device with 2 NVIDIA Tesla V100 (32G) GPUs. The detailed information of our experimental platform is shown in table 2.
Table 4: Ablation studies of (a) w/o SRA-I which use channel-independent strategy only, (b) w/o SRA-M which use channel-mixing strategy only, (c) w/o Bi which use forward direction Mamba block only, (d) w/o Residual that removes the residual connection, (e) S-Mamba and (f) PatchTST used for the benchmark models. The best results are in bold.
Models w/o SRA-I w/o SRA-M w/o Bi w/o Residual w/o M & w/ A S-Mamba PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 0.252 0.277 0.243 0.272 0.251 0.281 0.252 0.283 0.257 0.281 0.253 0.278 0.259 0.281
Traffic 0.453 0.287 0.404 0.272 0.435 0.289 0.439 0.306 0.417 0.281 0.415 0.276 0.481 0.305
Electricity 0.188 0.274 0.166 0.263 0.176 0.271 0.179 0.279 0.174 0.269 0.174 0.271 0.196 0.281
ETTh1 0.437 0.431 0.447 0.437 0.443 0.437 0.456 0.446 0.446 0.445 0.465 0.457 0.448 0.441
ETTh2 0.372 0.399 0.372 0.402 0.376 0.406 0.382 0.407 0.374 0.402 0.381 0.407 0.387 0.407
ETTm1 0.378 0.396 0.389 0.404 0.386 0.401 0.386 0.401 0.381 0.396 0.398 0.405 0.387 0.400
ETTm2 0.281 0.328 0.284 0.332 0.288 0.335 0.291 0.339 0.287 0.334 0.289 0.333 0.281 0.326
Solar 0.236 0.263 0.227 0.255 0.235 0.262 0.232 0.263 0.241 0.276 0.239 0.271 0.270 0.307
@@
Table 3 shows the results of long-term multivariate forecasting. Our model achieves superior performance compared with other baselines on all datasets with different predicting lengths. Compared with the current SOTA Transformer model iTransformer, Bi-Mamba+ reduces the MSE, MAE errors by 4.72% and 2.60% on average. The improvement of MSE and MAE comes to 3.76% and 2.67% compared with S-Mamba, which is also an SSM-based model using Mamba. It is worth noting that Bi-Mamba+ adopts channel-mixing strategy on Weather, Traffic, Electricity and Solar and channel-independent strategy on 4 ETT datasets. For channel-mixing strategy, the patching approach used by Bi-Mamba+ helps the model capture inter-series dependencies at a finer granularity compared with S-Mamba and iTransformer that transform the whole sequence into a token. For channel-independent strategy, the patch-wise tokens fed into Mamba blocks contain more semantic information, enhancing the model’s ability for capturing the inner evolutionary process of each univariate time series. Compared with PatchTST that adopt the same patching technique, our proposed Mamba+ block achieves more accurate predictions than self-attention.
@@
We verify the effectiveness of the SRA decider, the bidirectional module and the residual connection. Before we conduct the ablation experiments, we first export the results of the SRA decider to view the strategy choosing for different datasets. For Spearman correlation coefficient $\rho$, we typically use $\lambda>0.6$ to evaluate whether two variables have positive correlation strong enough to influence each other. The decision results under varying $\lambda$ are shown in Fig. LABEL:fig:decision_maker. For ETT datasets that have few variables, the model tends to adopt channel-independent strategy to emphasis intra-series dependencies, while for Weather, Traffic and Electricity that have large amount of variables, the model tends to adopt channel-mixing strategy to emphasis inter-series dependencies. We remove the SRA decider to form two specific Bi-Mamba+ models: (a) w/o SRA-I which only uses channel-independent tokenization strategy and (b) w/o SRA-M which only uses channel-mixing tokenization strategy. For the bidirectional Mamba+ encoder, we remove the backward direction Mamba+ to evaluate the effectiveness of bidirectional structure. This setting forms the following model: (c) w/o Bi. Besides, we remove the Add&Norm layer to evaluate the effect of residual connection: (d) w/o Residual. We also replace Mamba+ blocks with the dot-product Attention: (e) w/o M & w/ A.
We calculate the average MSE and MAE scores as shown in table 4. Without the SRA decider, the model utilizes either channel-mixing or channel independent tokenization strategies. The MSE and MAE increase by 0.95% and 0.77% for w/o SRA-M and 3.56% and 1.45% for w/o SRA-I respectively. Therefore, the proposed SRA decider makes the optimal choice of channel-independent and channel-mixing strategies and utilizing different tokenization strategies on different datasets produces better predicting performance.
When the backward Mamba block is removed, the predicting performance of MSE and MAE is deteriorated 3.25% and 2.44% respectively, which indicates that the design of bidirectional module enhances the model to get more comprehensive understanding of the MTS data.
The residual network enhances the expressive ability and feature utilization of the model by reusing early features. It can also reduce overfitting in the training process. The MSE and MAE of w/o Residual decrease by 4.37% and 4.05% respectively, indicating the effectiveness of the residual connection module. Full results of Ablation Study can be found in table 5 in Appendix.
It is worth noting that replacing Mamba+ with Attention causes performance degradation, indicating that Mamba may be a better choice for time series forecasting.
Figure 11: Results of varying patch length $P$ on Weather, Electricity, ETTh1 and ETTh2.
@@
To verify whether Bi-Mamba+ is sensitive to the hyper-parameters, we conduct experiments on (a) Spearman coefficient filter $\lambda$, (b) length of patches, (c) dropout (d) Conv-1D kernel size of Mamba+ and (e) state dimension size. For each setting, we freeze other parameters that are not currently fine-tuned and repeat the experiment 5 times with 40 epochs (with early-stop control). We record the average MSE of $H\in\{96,192,336,720\}$. Overall, different parameter settings do not cause significant changes in the results, which also shows the robustness of our proposed model.
Spearman coefficient filter $\lambda$.
The value $\lambda$ is used for filter out variable pairs that have positive correlations. A larger $\lambda$ corresponds to stronger positive correlations between two sequences. We set $\lambda\in\{0.2,0.4,0.6,0.8\}$ and calculate the decision ratio $r$ and decision indicator $ts$. The results are shown in Fig. LABEL:fig:decision_maker. Overall, different $\lambda$ does not frequently cause the change of tokenization strategy. For datasets that have large number of variables, such as Traffic and Electricity, the decision indicator $ts$ prefers to be equal to $1$, while for datasets with less variables such as ETTh1 and ETTh2, $ts$ is more likely to be 00. Note that the $ts$ may undergo a change from $\lambda=0.6$ to $\lambda=0.8$ on ETTh2 and ETTm2, resulting in different tokenization strategies. The results in table 4 indicate some correlations with the phenomenon. For ETTh2 and ETTm2, the tokenization strategy does not notably affect the predicting performance because of the existence of more highly correlated variables.
Patch length
We use varying patch length of $P\in\{\frac{1}{2}L,\frac{1}{4}L,\frac{1}{8}L,\frac{1}{12}L,\frac{1}{24}L\}$. We conduct experiments on Weather, Electricity, ETTh1 and ETTh2. For datasets that use channel-independent or channel-mixing tokenization strategy, different patch lengths reflect the intra- or inter-series dependencies at different timescales. The results are shown in Fig. 11. For Weather and ETT datasets, the model gets optimal performance with $P=\frac{1}{24}L$ or $P=\frac{1}{12}L$. These datasets have relatively less variables and the data distribution exhibit stronger non-stationarity locally, hence suitable for more fine-grained modeling. The optimal patch length for Electricity is $P=\frac{1}{4}L$, indicating that coarse-grained modeling is beneficial for processing the data that is overall stable. In general, varying lengths of patch does not cause significant performance differences, which indicates that our model is robust in capturing intra- and inter-series dependencies at different time scales.
Dropout
@@
The 1-D convolution module in Mamba is used for capturing temporal dependencies of the input series. We conduct experiments on different kernel size of 1, 2 and 3. The results in Fig. LABEL:fig:hyper(e) and Fig. LABEL:fig:hyper(f) show that Bi-Mamba+ gets optimal performance with the kernel size of 2. Overall, our model is not sensitive to the kernel size.
Figure 17: Results of varying look-back window size $L$ on Electricity and ETTm2. The selected baseline models are Autoformer, Transformer, PatchTST, Crossformer, DLinear and S-Mamba. Overall, Bi-Mamba+ gets more accurate prediction with the increased receptive field and shows better predicting performance than other models.
Long-term Dependencies Capturing
To evaluate the long-term dependencies capturing capability of Bi-Mamba+, we set $L\in\{48,96,192,336,720\}$ and record the average MSE scores of Bi-Mamba+, S-Mamba, Transformer, Autoformer, Crossformer and DLinear. We conduct experiments on Electricity and ETTm2. According to the SRA decider, these two datasets show preference to channel-mixing and channel-independent tokenization respectively, thus can more comprehensively reflect the adaptability of different models to the varying temporal evolutionary patterns.
Fig. 17 shows the results of the performance of models with varying look-back window size. Bi-Mamba+, S-Mamba, PatchTST and Crossformer get better predicting performance as the receptive field increases on Electricity. For previous channel-mixing models, Transformer does not show significant changes in predicting performance regardless of the size of the look-back window. Autoformer, however, even faces performance degradation at inflection point $L=96$. These two models may encounter the challenge of overfitting caused by point-wise tokens. We note that S-Mamba, PatchTST and S-Mamba do not achieve performance improvements similar to those on Electricity as the receptive fields increase on ETTm2. Compared to Electricity, the periodicity of ETTm2 is less obvious, making it more difficult to model the evolutionary patterns within the series. Previous channel-mixing models such as Autoformer encounters performance degradation at inflection point $L=96$, which may related to overfitting caused by point-wise tokens. Overall, Bi-Mamba+ can benefit from longer look-back window size and adapt to more complex temporal evolutionary patterns.
@@
We conduct the following experiments to comprehensively evaluate the model efficiency from (a) predicting accuracy, (b) memory usage and (c) training speed. We set $L=96,H=96$ as the forecasting task and use $Batch=32$ for ETTh1 and Traffic. The results are shown in Fig. LABEL:fig:efficiency(a) and Fig. LABEL:fig:efficiency(b). We freeze the patch length to $P=\frac{1}{4}L$ to provide consistent modeling process. We do not adopt Effcient Training(Liu et al. 2023) mechanism for Bi-Mamba+, S-Mamba and iTransformer on Traffic, in order to provide an intuitive comparison of the self-attention mechanism with Mamba.
For small datasets like ETTh1, Bi-Mamba+, S-Mamba, iTransformer, PatchTST and DLinear cost similar memory. For large datasets like Traffic, when the number of variables grows numerous, due to the the quadratic complexity of self-attention, the memory cost by iTransformer increases significantly, while Bi-Mamba+ and S-Mamba retain linear growth of resource usage, benefiting from the linear complexity of SSM. Therefore, Bi-Mamba+ strikes a good balance among predicting performance, training speed and memory usage. Overall, the computational cost of Mamba+ is less than self-attention mechanism in Transformer with the same embedding size $D$, which is reflected in the training speed and GPU memory usage. With the better predicting performance, faster training speed and less memory costs, Mamba-like models show great potential in LTSF.
Visualization
We visualize the long-term forecasting results of Bi-Mamba+, PatchTST, DLinear, and Autoformer in Fig. LABEL:fig:visual. We predict 96 steps on Traffic and Electricity. Bi-Mamba+ provides stable prediction results which is the closest to the ground truth.